OpenStudy (anonymous):

i give medals! find the inverse function of y=ln(x-4)

OpenStudy (anonymous):

write in equivalent exponential form as \[e^y=x-4\] then solve for \(x\) in one step

OpenStudy (anonymous):

Interchange the x and y variables, then solve for new y

OpenStudy (anonymous):

@satellite73, do you mean y?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then is it x=e^y + 4?

OpenStudy (anonymous):

you are trying to solve for \(x\) to find the inverse if you want to do it by interchanging \(x\) and \(y\) you get the same thing start with \[x=\ln(y-4)\] then write in exponential form as \[e^{x}=y-4\] and get \[y=e^x+4\]

OpenStudy (anonymous):

oh ok, thanks :)